I am writing a scientific notation problem the 34,500,000,000 and 72,500,000,000 my answer is 34.5x10^10 and 72.5x10^10 am I on the right track?
Answer by Alan3354(69443)

I am writing a scientific notation problem the 34,500,000,000 and 72,500,000,000 my answer is 34.5x10^10 and 72.5x10^10 am I on the right track?
------------------
You're one notch to the right of the track.
34,500,000,000 = 3.45*10^10
72,500,000,000 = 7.25*10^10
---------------
There's always one digit left of the decimal point.
Another notation is often used:
34,500,000,000 = 3.45*10^10 = 3.45E10
72,500,000,000 = 7.25*10^10 = 7.25E10
It's common in computer programming, and it's easier to type and say.